### v4

@@ 6 Columbia University
Optimization modeling plays a critical role in the application of Operations Research (OR) tools to address real-world problems, yet they pose challenges and require extensive expertise from OR experts. With the advent of large language models (LLMs), new opportunities have emerged to streamline and automate such task. However, current research predominantly relies on closed-source LLMs such as GPT-4, along with extensive prompt engineering techniques. This reliance stems from the scarcity of high-quality training datasets for optimization modeling, resulting in elevated costs, prolonged processing times, and privacy concerns. To address these challenges, our work is the first to propose a viable path for training open-source LLMs that are capable of optimization modeling and developing solver codes, eventually leading to a superior ability for automating optimization modeling and solving. Particularly, we introduce OR-Instruct, a semi-automated data synthesis framework for optimization modeling that enables customizable enhancements for specific scenarios or model types. We also introduce IndustryOR, the first industrial benchmark for evaluating LLMs in solving practical OR problems. We train several 7B-scale open-source LLMs using synthesized data (dubbed ORLMs), which exhibit significantly enhanced optimization modeling capabilities, achieving state-of-the-art performance across the NL4OPT, MAMO, and IndustryOR benchmarks. Additionally, our experiments highlight the potential of scaling law and reinforcement learning to further enhance the performance of ORLMs. The workflows and human-machine interaction paradigms of ORLMs in practical industrial applications are also discussed in the paper.
@@ 1 Introduction
The advent of large language models (LLMs) gives an emerging hope to the above challenge. As soon as its advent, these pre-trained LLMs have exhibited remarkable capabilities in rapid knowledge comprehension and cross-disciplinary learning, demonstrating remarkable success and transformative potential across various fields (Huang et al. 2024a), like medical consultation (Wang et al. 2023b), software development (Zaremba et al. 2021), offering stock market trading strategies (Chen et al. 2022b), and improving recommendation systems (Xu et al. 2024). In the realm of more abstract mathematical reasoning, research has increasingly incorporated more techniques to enhance the capabilities of large models, yielding promising results. For example, Trinh et al. (2024) train Alphaproof with hundreds of billions of synthetic data used to tackle complex olympiad geometry problems. Similarly, MathScale (Tang et al. 2024) offers a straightforward and scalable method to generate high-quality data for mathematical reasoning, establishing a foundation for further training of large language models in mathematical reasoning capabilities.
In recent years, with the advancements in pre-trained language models (PLMs) such as GPT-4 and Claude-3.5, a growing number of studies have explored automating the modeling and solving of optimization problems expressed in natural language. To produce complete solutions, including executable programs that leverage established solvers, existing approaches often employ prompt engineering techniques and multi-agent collaboration frameworks built on proprietary LLMs, such as Chain-of-Experts (Xiao et al. 2023), OptiMUS (AhmadiTeshnizi et al. 2024), and OptiGuide (Li et al. 2023a). However, utilizing online APIs for advanced large models offers no assurances of privacy protection, which is prohibitive in some sensitive industry applications. Additionally, these approaches are heavily reliant on the inherent capabilities of the foundational LLMs themselves, rendering them incapable of addressing the customized needs of users. As illustrated in Figure 1, even a basic optimization modeling problem presents challenges for the state-of-the-art foundational large language model, GPT-4. It demonstrates deficiencies in comprehending and transforming logical constraints, which imposes a limitation on methods dependent on these general closed-source LLMs.
@@
    Limited modeling capabilities. Contemporary advanced LLMs frequently exhibit errors in interpreting logical constraints or inadvertently incorporate nonlinear elements, complicating problem solution.
@@
To fill this gap, this paper introduces, for the first time, a new path for training an open-source large language model specifically designed for modeling and solving optimization problems. To address the core of the above challenges, we present a semi-automated framework for the synthesis of high-quality optimization modeling data, aimed at progressively enhancing the specialized modeling capabilities of LLMs. Our main contributions can be summarized as follows:
@@
    We utilize the synthetic data generated by OR-Instruct to train a series of open-source large language models, each approximately 7 billion parameters in size, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024). We designate these trained models as Operation Research Language Models (ORLMs) and observe a marked enhancement in their optimization modeling capabilities. Our experimental results show that our best-performing ORLM achieves state-of-the-art performance on the NL4OPT, MAMO, and IndustryOR benchmarks. Building upon this foundation, we conduct extensive numerical experiments to analyze the limitations of the ORLM and propose feasible solutions for its improvement. Our results highlight the promising prospects of ORLM in advancing the application and development of operations research, while also suggesting potential human-LLM collaboration frameworks. These frameworks demonstrate how ORLM could enhance the productivity of both industry and academia in the field of operations research.
@@
To our best knowledge, our work is the first initiative to train open-source LLMs specifically for optimization modeling in real-world industries. This approach significantly reduces the reliance on closed-source LLMs thus adding significant flexibility to the training process and greatly enhancing privacy preservation for industrial applications. Remarkably, our highest-performing ORLM achieves state-of-the-art results on the NL4OPT, MAMO, and IndustryOR benchmarks. This suggests that in the field of optimization modeling, our approach enables ORLMs with approximately 7 billion parameters to outperform many cutting-edge large language models, such as GPT-4, Llama-3.1, Qwen 2.5, and related agent-based frameworks.
@@
We also conduct a thorough analysis of the ORLM’s results, with numerous ablation experiments validating the effectiveness of our approach. Our findings show that, although ORLM performs relatively weak on some challenging datasets, several strategies can still ensure improvements in its performance. For instance, applying scaling laws to further increase the model size and data scale leads to continuous performance enhancements. Additionally, we find that ORLM has weak ranking capabilities for optimal optimization models. Using alignment strategies, such as reinforcement learning, also helps strengthen ORLM’s performance.
@@
We also want to comment that our proposed data synthesis method, the OR-Instruct, has the ability to facilitate customization to address particular needs of larger models, such as supply chain management, scheduling, inventory issues, or enhancing capabilities in linear programming, integer programming, and mixed integer programming. This synthetic data framework robustly addresses the scarcity of training data for optimization modeling, provides a valuable reference for advancing the application of operations research in various industrial sectors with the support of large models. In Section 6.1, we provide a detailed discussion of representative workflows driven by large models..
@@ 4.2 Model Finetuning and Inference
During training, we compute the loss only over the target completion. We employ the OR-Instruct dataset to train several open-source large language models (LLMs) of approximately 7 billion parameters, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024), which serve as our backbone models. We conduct a grid hyper-parameter search for each backbone and present the best-performing hyperparameters in Table 1. Prior to the final fine-tuning, we conducted a grid search with learning rates within {2e-5, 5e-5, 7e-5, 3e-6, 5e-6, 7e-6}, batch sizes within {64, 128, 256, 512}, and training epochs within {1, 2, 3} to determine the optimal settings. The models resulting from this process are referred to as ORLMs.
@@ 4.3 Evaluation and Baselines
In addition, we introduce the most well-known open-source large models for comparison, including Llama-3.1-lnstruct (Dubey et al. 2024), DeepSeek-v2-Chat (DeepSeek AI 2024), Qwen2-Instruct (Qwen Team 2024), and Mistral-Nemo (Mistral AI 2024). Each model is tasked with constructing optimization models and generating code to solve the test set problems.
@@
Our best-performing ORLMs, like ORLM-LLaMA-3 and ORLM-Qwen2.5, achieve state-of-the-art performance across four benchmarks with respect to GPT-4 and other open-source LLMs, surpassing Standard prompting based on GPT-4 by 42.2% in micro averageMicro average aggregates the outcomes of all classes to compute the metrics, heavily weighing the classes with more instances. This method is effective for evaluating the overall effectiveness of the model, particularly in datasets where some classes significantly outnumber others. and 55.4% in macro averageMacro average calculates individual metrics for each class without regard to class frequency, and then averages these metrics, treating all classes with equal importance. This approach highlights model performance on minority classes and is beneficial for assessing model fairness across diverse class distributions.. A comparison with human evaluation results reveals that ORLMs outperform senior undergraduate students on simpler problems (NL4OPT/MAMO-EasyLP) and approach expert-level performance. However, for more complex problems (IndustryOR/MAMO-ComplexLP), ORLMs still underperform relative to the average student, indicating that their problem-solving capabilities for complex tasks remain inadequate. We will analyze the underlying reasons for this in Section 5, highlighting that ORLMs still have the potential to surpass student performance and approach expert-level solutions on more challenging problems.
@@ 5.2 Ablation Study
We use linear programming as a case study, comparing ORLM’s performance when trained solely on linear programming data versus a mixed dataset including other problem types. Specifically, we construct two training sets, each comprising 7,049 instances: one set exclusively consisting of pure linear programming (LP) problems, and the other comprising a diverse mixture of problem typesThis dataset is comprised of 35% linear programming, 40% mixed integer programming, 25% other types.. We fine-tune the Meta-Llama3 8B model on both datasets, ensuring that hyper-parameter settings remain consistent across experiments. The test performance is summarized in Table 6:
@@
As illustrated in Table 6, even with identical test sizes, the inclusion of mixed problem types significantly enhances the performance of the ORLM-LLaMA-3-8B across various datasets. Notably, when examining the tests focused on linear programming, specifically MAMO-EasyLP and MAMO-ComplexLP, it is evident that integrating additional problem types is more effective in enhancing the ORLM’s performance on challenging linear programming problems. A plausible explanation for this phenomenon is that the incorporation of a broader variety of problems effectively strengthens the generalization ability of the ORLM-LLaMA-3-8B, allowing it to capture a wider modeling capacity and consequently improve modeling performance (See dong2023abilities).
@@ 6.2 Future Research Directions
7 Conclusion
In this paper, we propose a new path for the training of open-source large language models (LLMs) specifically tailored for optimization modeling. We delineate four critical requirements essential for the training dataset of optimization modeling LLMs and develop OR-Instruct, a semi-automated framework for generating synthetic data that meets these specific needs. Additionally, we introduce the IndustryOR benchmark, the first of its kind in the industry. We utilize the OR-Instruct data to train open-source LLMs with approximately 7 billion parameters, which significantly enhances their optimization modeling capabilities, achieving state-of-the-art performance across all test datasets.
In addition, this paper reveals the current limitations of ORLM, such as its overly simplistic output for complex problems, weak ability to rank optimal solutions, and poor learning capability from data. Our experimental results suggest that ORLM can be further enhanced by following scaling laws. For example, fine-tuning larger models with superior overall capabilities can enhance ORLM’s performance. Additionally, increasing the quantity and quality of training data or adjusting the data composition could bolster ORLM’s ability to address certain weaker problems. Furthermore, methods such as Reinforcement Learning from Human Feedback (RLHF), which train models based on preference data, could be introduced to further improve ORLM’s ranking ability for optimal solutions, potentially reaching expert-level performance. This finding offers significant insights for advancing ORLM’s applications in industry and highlights promising directions for future research.

### v5

@@ 6 Columbia University
Optimization modeling plays a critical role in the application of Operations Research (OR) tools to address real-world problems, yet they pose challenges and require extensive expertise from OR experts. With the advent of large language models (LLMs), new opportunities have emerged to streamline and automate such task. However, current research predominantly relies on closed-source LLMs such as GPT-4, along with extensive prompt engineering techniques. This reliance stems from the scarcity of high-quality training datasets for optimization modeling, resulting in elevated costs, prolonged processing times, and privacy concerns. To address these challenges, our work is the first to propose a viable path for training open-source LLMs that are capable of optimization modeling and developing solver codes, eventually leading to a superior ability for automating optimization modeling and solving. Particularly, we design the OR-Instruct, a semi-automated data synthesis framework for optimization modeling that enables customizable enhancements for specific scenarios or model types. This work also introduces IndustryOR, the first industrial benchmark for evaluating LLMs in solving practical OR problems. We train several 7B-scale open-source LLMs using synthesized data (dubbed ORLMs), which exhibit significantly enhanced optimization modeling capabilities, achieving competitive performance across the NL4OPT, MAMO, and IndustryOR benchmarks. Additionally, our experiments highlight the potential of scaling law and reinforcement learning to further enhance the performance of ORLMs. The workflows and human-machine interaction paradigms of ORLMs in practical industrial applications are also discussed in the paper.
@@ 1 Introduction
The advent of large language models (LLMs) gives an emerging hope to the above challenge. As soon as its advent, these pre-trained LLMs have exhibited remarkable capabilities in rapid knowledge comprehension and cross-disciplinary learning, demonstrating remarkable success and transformative potential across various fields (Huang et al. 2024a), like medical consultation (Wang et al. 2023b), software development (Zaremba et al. 2021), offering stock market trading strategies (Chen et al. 2022b), improving recommendation systems (Xu et al. 2024), and enhancing marketing research (Wang et al. 2024). In the realm of more abstract mathematical reasoning, research has increasingly incorporated more techniques to enhance the capabilities of large models, yielding promising results. For example, Trinh et al. (2024) train Alphaproof with hundreds of billions of synthetic data used to tackle complex olympiad geometry problems. Similarly, MathScale (Tang et al. 2024) offers a straightforward and scalable method to generate high-quality data for mathematical reasoning, establishing a foundation for further training of large language models in mathematical reasoning capabilities.
In recent years, with the advancements in pre-trained language models (PLMs) such as GPT-4 and Claude-3.5, a growing number of studies have explored automating the modeling and solving of optimization problems expressed in natural language. To produce complete solutions, including executable programs that leverage established solvers, existing approaches often employ prompt engineering techniques and multi-agent collaboration frameworks built on proprietary LLMs, such as Chain-of-Experts (Xiao et al. 2023), OptiMUS (AhmadiTeshnizi et al. 2024), and OptiGuide (Li et al. 2023a). However, utilizing online APIs for advanced large models offers no assurances of privacy protection, which is prohibitive in some sensitive industry applications. Additionally, these approaches are heavily reliant on the inherent capabilities of the foundational LLMs themselves, rendering them incapable of addressing the customized needs of users. As illustrated in Figure 1, even a basic optimization modeling problem presents challenges for the cutting-edge foundational large language model, GPT-4. It demonstrates deficiencies in comprehending and transforming logical constraints, which imposes a limitation on methods dependent on these general closed-source LLMs.
@@
    Limited modeling capabilities. Contemporary advanced LLMs frequently exhibit errors in interpreting logical constraints or inadvertently incorporate nonlinear elements, making it difficult to model real-world problems effectively.
@@
To fill this gap, this paper introduces, for the first time, a new path for training an open-source large language model specifically designed for modeling and solving optimization problems. To address the core of the above challenges, we present a semi-automated framework for the synthesis of high-quality optimization modeling data, aimed at progressively enhancing the specialized modeling capabilities of LLMs. We hope that this approach will inspire future research and serve as a foundation for the deeper integration of large models with operations research. Our main contributions can be summarized as follows:
@@
    We utilize the synthetic data generated by OR-Instruct to train a series of open-source large language models, each approximately 7 billion parameters in size, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024). We designate these trained models as Operation Research Language Models (ORLMs) and observe a marked enhancement in their optimization modeling capabilities. Our experimental results show that the best-performing ORLM achieves competitive performance on the NL4OPT, MAMO, and IndustryOR benchmarks. Building upon this foundation, we conduct extensive numerical experiments to analyze the limitations of the ORLM and propose feasible solutions for its improvement. The results highlight the promising prospects of ORLM in advancing the application and development of operations research, while also suggesting potential human-LLM collaboration frameworks. These frameworks demonstrate how ORLM could enhance the productivity of both industry and academia in the field of operations research.
@@
To our best knowledge, this work is the first initiative to train open-source LLMs specifically for optimization modeling in real-world industries. The proposed approach significantly reduces the reliance on closed-source LLMs thus adding significant flexibility to the training process and greatly enhancing privacy preservation for industrial applications. Remarkably, the highest-performing ORLM achieves competitive results on the NL4OPT, MAMO, and IndustryOR benchmarks. This suggests that in the field of optimization modeling, our approach enables ORLMs with approximately 7 billion parameters to outperform many cutting-edge large language models, such as GPT-4, Llama-3.1, Qwen 2.5, and related agent-based frameworks.
@@
We also conduct a thorough analysis of the ORLM’s results, with numerous ablation experiments validating the effectiveness of our approach. The results show that, although ORLM performs relatively weak on some challenging datasets, several strategies can still ensure improvements in its performance. For instance, applying scaling laws to further increase the model size and data scale leads to continuous performance enhancements. Additionally, we find that ORLM has weak ranking capabilities for optimal optimization models. Using alignment strategies, such as reinforcement learning, also helps strengthen ORLM’s performance.
@@
Furthermore, we want to comment that the proposed data synthesis method, the OR-Instruct, has the ability to facilitate customization to address particular needs of larger models, such as supply chain management, scheduling, inventory issues, or enhancing capabilities in linear programming, integer programming, and mixed integer programming. This synthetic data framework robustly addresses the scarcity of training data for optimization modeling, provides a valuable reference for advancing the application of operations research in various industrial sectors with the support of large models. In Section 6.1, we provide a detailed discussion of representative workflows driven by large models.
@@ 4.2 Model Finetuning and Inference
During training, we compute the loss only over the target completion. We employ the OR-Instruct dataset to train several open-source large language models (LLMs) of approximately 7 billion parameters, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024), which serve as our backbone models. We conduct a grid hyper-parameter search for each backbone and present the best-performing hyperparameters in Table 1.  [^1] Prior to the final fine-tuning, we conducted a grid search with learning rates within {2e-5, 5e-5, 7e-5, 3e-6, 5e-6, 7e-6}, batch sizes within {64, 128, 256, 512}, and training epochs within {1, 2, 3} to determine the optimal settings. The models resulting from this process are referred to as ORLMs.
@@ 4.3 Evaluation and Baselines
In addition, we introduce the most well-known open-source large models for comparison, including Llama-3.1-lnstruct (Dubey et al. 2024), DeepSeek-v2-Chat (DeepSeek AI 2024), DeepSeek-R1 (Guo et al. 2025), Qwen2-Instruct (Qwen Team 2024), and Mistral-Nemo (Mistral AI 2024). Each model is tasked with constructing optimization models and generating code to solve the test set problems.
@@
Our best-performing ORLMs, like ORLM-LLaMA-3 and ORLM-Qwen2.5, achieve state-of-the-art performance across four benchmarks with respect to GPT-4 and other open-source LLMs [^2] Due to the severe scarcity of API resources, we are unable to conduct large-scale modeling tests on many state-of-the-art LLMs, such as DeepSeek-R1 (671B)., surpassing Standard prompting based on GPT-4 by 42.2% in micro average [^3] Micro average aggregates the outcomes of all classes to compute the metrics, heavily weighing the classes with more instances. This method is effective for evaluating the overall effectiveness of the model, particularly in datasets where some classes significantly outnumber others. and 55.4% in macro average [^4] Macro average calculates individual metrics for each class without regard to class frequency, and then averages these metrics, treating all classes with equal importance. This approach highlights model performance on minority classes and is beneficial for assessing model fairness across diverse class distributions. A comparison with human evaluation results reveals that ORLMs outperform senior undergraduate students on simpler problems (NL4OPT/MAMO-EasyLP) and approach expert-level performance. However, for more complex problems (IndustryOR/MAMO-ComplexLP), ORLMs still underperform relative to the average student, indicating that their problem-solving capabilities for complex tasks remain inadequate. We will analyze the underlying reasons for this in Section 5, highlighting that ORLMs still have the potential to surpass student performance and approach expert-level solutions on more challenging problems.
@@
A comparative study on problem-solving efficiency. To verify the effectiveness of ORLM in human–AI collaboration, we conduct a comparative experiment to demonstrate its impact on algorithm engineers in solving practical problems. Specifically, we recruit a group of 30 people, consisting of experts and senior undergraduate students, dividing them evenly into two groups, A and B. Each group comprises 7 experts and 8 undergraduate students, with comparable proficiency levels. We design 7 problems with varying difficulty levels, ranging from textbook exercises to those approximating industrial complexity. Group A is tasked with independently modeling and programming the solutions without the aid of ORLM, while Group B utilizes ORLM for problem-solving and programming. We assess the effectiveness of ORLM based on two dimensions: accuracy and solution time. The results are presented in Figure 8. We acknowledge that the sample size of the experiment is limited due to the difficulty of finding suitable subjects; nevertheless, we perform statistical analysis on the limited data and find that the comparison is still statistically significant. In the following, we detail our statistical analysis.
[Comparison of solution time] Refer to caption [Comparison of accuracy] Refer to caption
Figure 8: A comparative study on problem-solving efficiency
Table 4 shows the results of our statistical analysis. Specifically, for each metric (difference in accuracy for total/experts/students, difference in time for total/experts/students), we first select an appropriate statistical test based on the distributional properties of the data, as indicated in the third column of the table. For solution time of each subgroup, we find that the distribution successfully passes the normality test and meets the assumption of homogeneity of variance. Consequently, we employ a standard t-test for statistical analysis. In contrast, for accuracy, the distribution does not conform to normality; therefore, we utilize the Mann-Whitney U test as a non-parametric alternative (see Hogg et al. 2019 for reference). Subsequently, we compute the test statistic and the corresponding $p$-value for each test, which are reported in the 4th and 5th columns, respectively. We also calculate the 95% confidence interval for the difference between the two groups based on each test and present the results in the 6th column.
To further demonstrate the robustness of our results, the last two columns provide additional analyses, including the effect size (Hedge’s g) and the statistical power (Cohen 2013). Notably, Hedge’s g is independent of sample size, making it a more objective measure of the generalizability and strength of the statistical significance. Meanwhile, statistical power serves as an indicator of the reliability and accuracy of our results. The detailed testing procedures and statistical indicators are provided in Appendix 8.6.
Table 4: Summary of tests and effect sizes
Type Test Method Statistic Testing
Stat $p$-value Confidence Interval Hedge’s g Stat Power
Total Accuracy Mann-Whitney U test 192.00 0.001 (0.14, 0.24) 1.46 0.97
Time Mann-Whitney U test 0.00 0.000 (-134.02, -112.35) -4.43 1.00
Expert Accuracy Mann-Whitney U test 44.00 0.008 (0.10, 0.25) 1.81 0.87
Time t-test -22.50 0.000 (-117.61, -96.85) -11.26 1.00
Student Accuracy Mann-Whitney U test 56.5 0.006 (0.14, 0.26) 1.56 0.82
Time t-test -20.27 0.000 (-151.29, -122.34) -9.58 1.00
  • a)
    For the criteria used in selecting between the Mann-Whitney U test and the independent samples t-test, please refer to Appendix 8.6.
For Table 4, we can see that at a 95% confidence level, the use of ORLM significantly improves both solution time and accuracy across all populations (total, expert, and student samples). The confidence intervals suggest that, after implementing ORLM, accuracy improves by approximately 10%–25%. Additionally, the total time saved in solving optimization modeling problems is 1.8–2.2 hours, with experts saving approximately 1.5–2 hours and students saving around 2–2.5 hours. The absolute value of Hedge’s g $>1$ indicates a substantial effect size, highlighting a large difference between experimental and control groups. Additionally, statistical power $>0.8$, ensuring a high probability of detecting a true effect and minimizing Type II error risk.
In addition, we collect feedback from engineers on their experience using ORLM, as shown in Table 5. None of the engineers rate ORLM as "useless" or "minimal help". Instead, the majority provide highly positive evaluations ("significant help" or "highly effective"), indicating that ORLM is both efficient and effective in assisting their work. These demonstrate the significant potential of ORLM to enhance the application of operations research in industry through human–AI collaboration. We will further discuss the application value of ORLM in Section 6.1.
Table 5: Feedback from participants on ORLM effectiveness
Category Useless Minimal Help Moderate Help Significant Help Highly Effective
Percentage (%) 0.0 0.0 13.3 60.0 26.7
@@ 5.2 Ablation Study
We use linear programming as a case study, comparing ORLM’s performance when trained solely on linear programming data versus a mixed dataset including other problem types. Specifically, we construct two training sets, each comprising 7,049 instances: one set exclusively consisting of pure linear programming (LP) problems, and the other comprising a diverse mixture of problem types [^5] This dataset is comprised of 35% linear programming, 40% mixed integer programming, 25% other types. We fine-tune the Meta-Llama3 8B model on both datasets, ensuring that hyper-parameter settings remain consistent across experiments. The test performance is summarized in Table 8:
@@
As illustrated in Table 8, even with identical test sizes, the inclusion of mixed problem types significantly enhances the performance of the ORLM-LLaMA-3-8B across various datasets. Notably, when examining the tests focused on linear programming, specifically MAMO-EasyLP and MAMO-ComplexLP, it is evident that integrating additional problem types is more effective in enhancing the ORLM’s performance on challenging linear programming problems. A plausible explanation for this phenomenon is that the incorporation of a broader variety of problems effectively strengthens the generalization ability of the ORLM-LLaMA-3-8B, allowing it to capture a wider modeling capacity and consequently improve modeling performance (See Dong et al. 2023).
@@ 6.2 Future Research Directions
7 Concluding Remarks
In this paper, we propose a new path for the training of open-source large language models (LLMs) specifically tailored for optimization modeling. We characterize four critical requirements essential for the training dataset of optimization modeling LLMs and develop OR-Instruct, a semi-automated framework for generating synthetic data that meets these specific needs. Additionally, we introduce the IndustryOR benchmark, the first of its kind in the industry. We utilize the OR-Instruct data to train open-source LLMs with approximately 7 billion parameters, which significantly enhances their optimization modeling capabilities, achieving competitive performance across all test datasets.
This paper also identifies the current limitations of ORLM, such as its overly simplistic output for complex problems, weak ability to rank optimal solutions, and poor learning capability from data. Our experiments suggest that ORLM can be further enhanced by leveraging scaling laws or incorporating methods such as reinforcement learning (RL), potentially reaching expert-level performance. We hope these findings provide valuable insights for the future advancement of ORLM’s applications in industry and open promising avenues for subsequent research.
Finally, we note that during the revision process of this paper, numerous high-performing large models have emerged, such as DeepSeek-R1 and Grok-3. While these new models demonstrate impressive capabilities, further enhancing their performance in optimization modeling requires additional refinement. The framework proposed in this paper offers a valuable reference for such improvements, including synthesizing high-quality data to strengthen modeling comprehension, or considering reinforcement learning strategies to align ranking capabilities. By leveraging our framework as a promising starting point, the inherent strengths of large models can be more effectively translated into optimization modeling capabilities.